Grade 11General Physics

relation between time and distrance is
t=ax^2+bx where a and b are constants.
retardation is ------
options are
a-2av^3
b-2bv^3
c-2abv^3
d-2b^2 v^3

aditya
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
t=ax^2+bx
differtiate it:
 
1 = 2axdx/dt + bdx/dt
dx/dt = 1/(2ax + b)
v = 1/(2ax + b)
Now, acc. = vdv/dx
 
dv/dx = -2a/(2ax +b)^2
 
acc. = -2a/(2ax + b)^3
Or
acc. = -2av^3